### v4

Abstract
Phase-space versions of quantum mechanics – from Wigner’s original distribution to modern discrete-qudit constructions – represent some states with negative quasi-probabilities. Conventional Shannon and Rényi entropies become complex-valued in this setting and lose their operational meaning. Building on the axiomatic treatments of Rényi (1961) and Daróczy (1963), we develop a conservative extension that applies to signed finite phase spaces and identify a single admissible entropy family, which we call signed Rényi $\alpha$-entropy (for a free parameter $\alpha>0$). The obvious signed Shannon candidate is ruled out because it violates extensivity. We prove four results that bolster the usefulness of the new measure. (i) It serves as a witness of negative probability. (ii) For $\alpha>1$, it is Schur-concave, delivering the intuitive property that mixing increases entropy. (iii) The same parametric family obeys a quantum H-theorem, namely, that under de-phasing dynamics entropy cannot decrease. (iv) The $2$-entropy is conserved under discrete Moyal-bracket dynamics, mirroring conservation of von Neumann entropy under unitary evolution on Hilbert space. We also comment on interpreting the Rényi order parameter as an inverse temperature. Overall, we believe that our investigation provides good evidence that our axiomatically derived signed Rényi entropy may be a useful addition to existing entropy measures employed in quantum information, foundations, and thermodynamics.
1 Introduction
Phase-space formulations of quantum mechanics – beginning with Wigner’s quasi-probability distribution (Wigner, 1932) and including Feynman’s comments on negative probability (Feynman, 1987) and the discrete constructions for qudits (Wootters, 1987; Gibbons, Hoffman, and Wootters, 2004) – assign negative weights to certain states. When a logarithm operatives on these negative quantities, the familiar Shannon or Rényi entropies become complex-valued and lose their operational meaning in terms of the amount of uncertainty about a system. This poses an important foundational question:
What axioms must an entropy satisfy to remain real-valued and physically informative when its input can be negative?
A clue to the answer can be found in the Rényi (1961) axiomatization of entropy that generalized Shannon entropy (Shannon, 1948) and which Rényi later used in various problems in probability theory and information theory. (Applications to ergodicity of Markov chains and the central limit theorem can be found in Rényi, 1970, pp.598-603.) Since then, Rényi entropy has been used in many fields (Csiszár, 2008). Two instances are quantum information and quantum thermodynamics, which are the primary motivation for the current paper. See Coles et al. (2017) and Goold et al. (2016) for comprehensive surveys.
The usual definition of entropy in quantum mechanics is von Neumann entropy (von Neumann, 1932), which is the natural analog to Shannon entropy. One can also define the Hilbert-space analog to Rényi entropy, as in the works above. Our goal is to find a natural definition on phase space.
In this paper, we return to the Rényi-Daróczy axioms (Rényi, 1961; Daróczy, 1963) and modify them so that entropy retains its natural meaning in the presence of negative measure. We obtain a characterization of what we called signed Rényi $\alpha$-entropy:
$H^{\pm}_{\alpha}(P)=-\frac{1}{\alpha-1}\log_{2}\bigl{[}\frac{\sum_{i=1}^{n}|p_{i}|^{\alpha}}{|\sum_{i=1}^{n}p_{i}|}\bigr{]},$ (1)
where $P=(p_{1},\ldots,p_{n})$ is a (real-valued) signed measure on phase space and $\alpha>0$ with $\alpha\not=1$ is a free parameter. Clearly, Equation (1) reduces to the usual formula for Rényi $\alpha$-entropy when $P$ is an ordinary probability measure.
An immediate question is why we do not adopt the obvious signed analog to Shannon entropy:
$H^{\pm}_{1}(P)=-\sum_{i=1}^{n}|p_{i}|\log_{2}|p_{i}|.$ (2)
The answer is that this formula fails the basic physical requirement of extensivity: $H^{\pm}_{1}(P*Q)\not=H^{\pm}_{1}(P)+H^{\pm}_{1}(Q)$. This requirement seems basic for both information and thermodynamic interpretations of entropy. (Notice that, as must then be the case, $H^{\pm}_{\alpha}(P)$ does not yield signed Shannon entropy in the $\alpha\rightarrow 1$ limit. In fact, it will typically diverge at $\alpha=1$.)
We further justify our choice of entropy for signed phase space by showing that:
Negativity Witness $H^{\pm}_{\alpha}(P)$ witnesses probabilistic negativity via a simple sign test;
Schur-Concavity $H^{\pm}_{\alpha}(P)$ is Schur-concave for all $\alpha>1$;
Quantum H-Theorem $H^{\pm}_{\alpha}(P)$ is non-decreasing, if $\alpha>1$, under a de-phasing assumption on dynamics;
Unitarity Analog $H^{\pm}_{2}(P)$ is preserved under discrete Moyal-bracket evolution, paralleling preservation of Hilbert-space entropies under unitary evolution.
Signed Rényi entropies already underpin an entropic characterization of the qubit via an uncertainty principle (Brandenburger, La Mura, and Zoble, 2022) and feature in recent analyses of quantum and superquantum correlations (e.g., Aw et al., 2023; Onggadinata, Kurzynski, and Kaszlikowski, 2023a, 2023b; Onggadinata et al., 2024). The links between Hilbert-space Rényi entropy and thermodynamics, many of which are surveyed in Goold et al., 2016), have been explored in detail by Baez (2022), Bakiev, Nakshidze, and Savchenko (2020); Jizba and Arimitsu (2004); Lu and Grover (2019); and Stéphan, Misguich, and Pasquier (2010). A recent Rényi-like measure of interest is the $\alpha$-logarithmic negativity (Wang and Wilde, 2020). Ji et al. (2024) connect relative Rényi entropy to the quantum state exclusion task. By supplying an axiomatic foundation for Rényi entropy on phase space and establishing some preliminary properties, we hope to point the way to other uses of phase-space Rényi entropy in quantum information, foundations, and thermodynamics.
In Section 2, we review the classic Rényi (1961) axioms for ordinary measures. In Section 3, we introduce our modification of the axiom set to encompass signed measures and state our main characterization theorem. In Sections 4 and 5, we show how our signed Rényi entropy witnesses negativity and satisfies Schur concavity, and we give examples to show that the signed Shannon entropy candidate fails both properties. Section 6 uses our formula to state and prove an abstract-level quantum H-theorem. Section 7 establishes the baseline that our signed Rényi entropy is conserved under Moyal-bracket evolution on phase space, which is a necessary check that we have identified a well-behaved analog to von Neumann (or Rényi) entropy on Hilbert space. Section 8 examines a renormalized version of our formula and looks briefly at interpreting the $\alpha$ parameter as an inverse temperature. Section 9 concludes.
2 Axioms for Rényi Entropy
We follow the approach in Rényi (1961) by first axiomatizing entropy for the class of non-negative measures with total weight less than or equal to $1$ and then specializing to probability measures. Fix a finite set of states $X=\{x_{1},\ldots,x_{n}\}$. A generalized probability measure on $X$ is a tuple $P=(p_{1},\ldots,p_{n})$ where each $p_{i}\in\mathbb{R}_{+}$ and $\sum_{i=1}^{n}p_{i}\leq 1$. The quantity $w(P)=\sum_{i=1}^{n}p_{i}$ is called the weight of $P$. We assume $w(P)>0$.
Given two generalized probability measures $P=(p_{1},\ldots,p_{m})$ and $Q=(q_{1},\ldots,q_{n})$, we denote by $P*Q$ the generalized probability measure which is the direct product:
$(p_{1}q_{1},\ldots,p_{1}q_{n},\ldots,p_{m}q_{1},\ldots,p_{m}q_{n}).$ (3)
Also, we denote by $P\cup Q$ the generalized probability measure which is the (direct) sum:
$(p_{1},\ldots,p_{m},q_{1},\ldots,q_{n}),$ (4)
whenever it is well-defined, i.e., $\sum_{i}p_{i}+\sum_{j}q_{j}\leq 1$. Finally, we write $(p)$ for the generalized probability measure consisting of the single real number $p>0$.
The Rényi (1961) axioms are:
Axiom 1.
(Symmetry)    $H(P)$ is a symmetric function of the elements of $P$.
Axiom 2.
(Continuity)    $H((p))$ is continuous for all $0<p\leq 1$.
Axiom 3.
(Calibration)    $H((\frac{1}{2}))=1$.
Axiom 4.
(Extensivity)    $H(P*Q)=H(P)+H(Q)$.
Axiom 5.
(Mean-Value Property)    There is a strictly monotone and continuous function $g:\mathbb{R}\rightarrow\mathbb{R}$ such that for any $P,Q$, whenever $H(P\cup Q)$ is well-defined:
$H(P\cup Q)=g^{-1}\big{[}\frac{w(P)g(H(P))+w(Q)g(H(Q))}{w(P\cup Q)}\big{]}.$ (5)
Rényi (1961, Theorem 2) proved that if the function $g$ in Axiom 5 takes the form $g(x)=-dx+e$, for constants $d\not=0$ and $e$, then Axioms 1-5 characterize the entropy functional:
${}H_{1}(P)=-\frac{\sum_{i=1}^{n}p_{i}\log_{2}p_{i}}{\sum_{i=1}^{n}p_{i}},$ (6)
while, if $g(x)=-d2^{(1-\alpha)x}+e$ for $\alpha>0$ with $\alpha\not=1$, then Axioms 1-5 characterize the family of entropy functionals:
${}H_{\alpha}(P)=-\frac{1}{\alpha-1}\log_{2}\big{[}\frac{\sum_{i=1}^{n}p_{i}^{\alpha}}{\sum_{i=1}^{n}p_{i}}\big{]}.$ (7)
The entropies in Equations 6 and 7 are the familiar Shannon and Rényi entropies, defined for generalized probability measures. As is well known, they can be related via an application of l’Hôpital’s rule:
${}\lim_{\alpha\rightarrow 1}H_{\alpha}(Q)=H_{1}(Q).$ (8)
Daróczy (1963) proved that Axioms 1-5 are, in fact, fully characterized by the two functional forms $g(x)=-dx+e$ and $g(x)=-d2^{(1-\alpha)x}+e$ with the preceding parameter restrictions on $d,e$, and $\alpha$. (He also observed that Axiom 1 is not needed, because it follows from Axiom 5.)
3 Extension to Signed Measures
A signed measure $P$ on $X$ is a tuple $P=(p_{1},\ldots,p_{n})$ where each $p_{i}\in\mathbb{R}$. Thus, the components $p_{i}$ are no longer required to be non-negative. We impose $w(P)\neq 0$ but not $w(P)=1$ (except when $P$ is a signed probability measure). The notations $P*Q$ and $P\cup Q$ have the same meanings as in the previous section and the latter is well-defined when $\sum_{i}p_{i}+\sum_{j}q_{j}\neq 0$. We now state our modified set of axioms.
Axiom 0. (Real-Valuedness) $H(P)$ is a non-constant real-valued function of $P$.
Axiom 2. (Continuity) $H((p))$ is continuous for all $p\neq 0$.
Axiom 3. (Calibration) $H\left(\left(\frac{1}{2}\right)\right)=1$.
Axiom 4. (Extensivity) $H(P*Q)=H(P)+H(Q)$.
Axiom 5. (Mean-Value Property) There is a strictly monotone and continuous function $g:\mathbb{R}\to\mathbb{R}$ such that for any $P,Q$, whenever $H(P\cup Q)$ is well-defined:
$H(P\cup Q)=g^{-1}\bigl{[}\frac{w(P)g(H(P))+w(Q)g(H(Q))}{w(P\cup Q)}\bigr{]}.$ (9)
Some comments on the axioms. Axiom 0 ensures that entropy, as a measure of uncertainty or lack of information, is real-valued, which is critical for signed measures where standard entropy formulas may yield complex values due to negative components. Axiom 2 extends Axiom 2 to signed measures, excluding $p=0$. Axioms 3, 4, and 5 are as in Rényi (1961), stated for signed measures.
Theorem 1.
Axioms 0, 2, 3, 4, and 5 hold if and only if $H(P)$ is given by:
$H^{\pm}_{\alpha}(P)=-\frac{1}{\alpha-1}\log_{2}\bigl{[}\frac{\sum_{i=1}^{n}|p_{i}|^{\alpha}}{\left|\sum_{i=1}^{n}p_{i}\right|}\bigr{]},$ (10)
where $\alpha\in\mathbb{R}$ is a free parameter with $\alpha>0$ and $\alpha\neq 1$.
The proof of Theorem 1 is in the Appendix. Here, we make some observations. An obvious signed analog to ordinary Shannon entropy is:
$H^{\pm}_{1}(P)=-\frac{\sum_{i=1}^{n}|p_{i}|\log_{2}|p_{i}|}{|\sum_{i=1}^{n}p_{i}|},$ (11)
defined when $w(P)\not=0$. However, this formula violates the Extensivity axiom. (From now on, all logs unless otherwise noted will be to base $2$.)
Example 1.
Let $P$ and $Q$ be the same signed probability measure $(2,-1)$. Then:
$H^{\pm}_{1}(P)=H^{\pm}_{1}(Q)=-(2\log 2+1\log 1)=-2.$ (12)
For the product measure $P*Q=(4,-2,-2,1)$:
$H^{\pm}_{1}(P*Q)=-(4\log 4+2\cdot 2\log 2+1\log 1)=-12\not=-4=H^{\pm}_{1}(P)+H^{\pm}_{1}(Q),$ (13)
showing that Axiom 4 does not hold. By contrast, for signed Rényi $\alpha$-entropy:
$H^{\pm}_{\alpha}(P)=H^{\pm}_{\alpha}(Q)=-\frac{1}{\alpha-1}\log(2^{\alpha}+1),$ (14)
and:
$H^{\pm}_{\alpha}(P*Q)=-\frac{1}{\alpha-1}\log(4^{\alpha}+2\cdot 2^{\alpha}+1)=-\frac{1}{\alpha-1}\log\bigl{[}(2^{\alpha}+1)^{2}\bigr{]},$ (15)
consistent with Axiom 4.
We see the failure of signed Shannon entropy to obey Extensivity as significant evidence against its value as an entropy measure – and as strong evidence in favor of the value of signed Rényi entropy. Nevertheless, we will hold signed Shannon entropy in mind as we explore various properties of signed Rényi entropy, so that we can compare their performance more broadly.
Note from Example 1 that signed Shannon entropy (Equation 11) cannot arise as the $\alpha\rightarrow 1$ limit of signed Rényi entropy $H^{\pm}_{\alpha}(P)$. Indeed, it can be seen from Equation 10 that signed Rényi entropy will generally diverge at $\alpha=1$. We do not see this as a shortcoming of our formula, which explicitly rules out $\alpha=1$. (But see Section 8.)
A closely related paper to this one is Koukouledikis and Jennings (2022). These authors do not take an axiomatic approach, but, instead, directly assume that $\alpha=2a/(2b-1)$, where $a,b$ are positive integers with $a\geq b$, so that $p_{i}^{\alpha}$ is non-negative real-valued. Our entropy formula is defined for all $\alpha>0$, but it coincides with the Koukouledikis-Jennings one when $\alpha$ is restricted as in their paper. We also treat unnormalized measures. In the next two sections, we parallel Koukouledikis and Jennings (2022) in examining how non-classicality of probabilities is witnessed and in examining Schur-concavity properties.
4 Witnessing Negativity
In this section, we show that signed Rényi entropy as just defined witnesses non-classicality of measures. This result mirrors closely and also extends the domain of Koukouledikis and Jennings (2020, Theorem 10). As we will see, signed Shannon entropy does not witness non-classicality – at least, not in this way.
Theorem 2.
A (signed) measure $P=(p_{1},\ldots,p_{n})$ contains at least one strictly negative component $p_{i}$ if and only if there is a value of the free parameter $\alpha>1$ such that the Rényi $\alpha$-entropy $H^{\pm}_{\alpha}(P)$ is strictly negative.
Proof.
Suppose $p_{i}<0$ for some $i=1,\ldots,n$. Then $\sum_{i=1}^{n}|p_{i}|>|\sum_{i=1}^{n}p_{i}|$. It follows that we can find positive numbers $\epsilon$ and $r$ such that:
$\frac{\sum_{i=1}^{n}|p_{i}|^{1+\epsilon}}{|\sum_{i=1}^{n}p_{i}|}=r>1,$ (16)
But then $H^{\pm}_{1+\epsilon}(P)=-1/\epsilon\times\log_{2}(r)$, which is less than 00.
For the converse, consider a general (unnormalized) non-negative measure $P$. We can copy the usual argument made for non-negativity of Rényi entropy in the case of probability measures. If $10<\alpha<$ 1, then $\sum_{i=1}^{n}p_{i}^{\alpha}\geq\sum_{i=1}^{n}p_{i}$, and, if $\alpha>1$, then $\sum_{i=1}^{n}p_{i}^{\alpha}\leq\sum_{i=1}^{n}p_{i}$, so that in both cases $H^{\pm}_{\alpha}(P)\geq 0$. ∎
Observe that Theorem 2 works by quantifying over $\alpha$. This is natural since Rényi entropy, whether signed or unsigned, is a family of functionals, not a single functional, and so it makes sense that a test – such as this one for negativity of probability – quantifies over the free parameter. The next example shows that signed Shannon entropy does not witness negativity the same way.
Example 2.
Consider the signed probability measure $P=(2,-\frac{1}{n},\ldots,-\frac{1}{n})$ where the term $-\frac{1}{n}$ is repeated $n$ times. We can calculate:
${}H^{\pm}_{1}(P)=-2\log 2-n\frac{1}{n}\log(\frac{1}{n})=\log n-2,$ (17)
which is strictly positive as long as $n>4$. To check that signed Rényi entropy acts as a witness in this example, we also calculate:
${}H^{\pm}_{1+\epsilon}(P)=-\frac{1}{\epsilon}\log\bigl{[}2^{1+\epsilon}+n\times(\frac{1}{n})^{1+\epsilon}\bigr{]}\approx-\frac{1}{\epsilon}\log 3,$ (18)
for small positive $\epsilon$, which is less than 00 as required by Theorem 2.
5 Schur-Concavity
We now turn to majorization, that is, the notion that one measure is “more uniform” than another one. This is a basic test for any candidate for an entropy measure. Recall that a function $f:\cal D\rightarrow\mathbb{R}$, where ${\cal D}\subseteq\mathbb{R}^{n}$, is Schur-concave on $\cal D$ if $f(P)\geq f(Q)$ whenever $P\in\cal D$ is majorized by $Q\in\cal D$, that is:
$\sum_{i=1}^{m}p_{[i]}$ $\leq\sum_{i=1}^{m}q_{[i]}\,\,\text{for all}\,\,m\leq n-1,~{}$ (19)
$\sum_{i=1}^{n}p_{[i]}$ $=\sum_{i=1}^{n}q_{[i]},~{}$ (20)
where $(p_{[1]},p_{[2]},\ldots,p_{[n]})$ is a decreasing rearrangement of $P$, and similarly for $Q$.
Theorem 3.
If $\alpha>1$, then signed Rényi entropy $H^{\pm}_{\alpha}$ is Schur-concave on $\{P\in\mathbb{R}^{n}:\sum_{i}p_{i}\not=0\}$.
Proof.
We first consider $P,Q\in\mathbb{R}^{n}$ without imposing the restrictions $\sum_{i}p_{i}\not=0$ and $\sum_{i}q_{i}\not=0$. Since $|p_{i}|^{\alpha}$ is convex for all $\alpha>1$, it follows from Marshall, Olkin, and Arnold (2011, Proposition C.1, p.92) that $\sum_{i}|p_{i}|^{\alpha}$ is Schur-convex on $\mathbb{R}^{n}$. Therefore, if $P$ is majorized by $Q$:
${}\sum_{i=1}^{n}|p_{i}|^{\alpha}\leq\sum_{i=1}^{n}|q_{i}|^{\alpha}.$ (21)
Now assume $\sum_{i}p_{i}=\sum_{i}q_{i}=w\not=0$. The map $x\mapsto-1/(\alpha-1)\times\log(x/w)$ is strictly decreasing for $\alpha>1$. Applying it to both sides of Equation 21 reverses the inequality, yielding $H^{\pm}_{\alpha}(P)\geq H^{\pm}_{\alpha}(Q)$. This establishes that $H^{\pm}_{\alpha}$ is Schur-concave on $\{P\in\mathbb{R}^{n}:\sum_{i}p_{i}=w\not=0\}$. ∎
Koukouledikis and Jennings (2022, Theorem 9) state that Rényi entropy for the parametric family $\alpha=2a/(2b-1)$, where $a,b$ are positive integers with $a\geq b$, is Schur-concave for signed probability measures. Our Theorem 3 extends the domain of their result.
Example 3.
Consider the signed probability measures $P=(0.08,0.45,0.47)$ and $Q=(-0.3,0.6,0.7)$. Then $P$ is majorized by $Q$. The corresponding signed Shannon entropies are:
$H^{\pm}_{1}(P)$ $=-(0.08\log 0.08+0.45\log 0.45+0.47\log 0.47)\approx 1.3219,$ (22)
$H^{\pm}_{1}(Q)$ $=-(0.3\log 0.3+0.6\log 0.6+0.7\log 0.7)\approx 1.3235,$ (23)
so that $H^{\pm}_{1}(P)<H^{\pm}_{1}(Q)$, contradicting Schur-concavity. By contrast, we find the signed Rényi $2$-entropies satisfy $H^{\pm}_{2}(P)\approx 1.2183$ and $H^{\pm}_{2}(Q)\approx 0.0893$, in line with Theorem 3.
6 A Quantum H-Theorem
We now make use of our concept of signed Rényi entropy is to obtain a quantum H-theorem. The puzzle of what, given unitarity, is the appropriate quantum analog to the classical Boltzmann H-theorem has a long history. Pauli (1928) developed a time-irreversible quantum master equation by assuming that off-diagonal elements of the density matrix vanish over time owing to random-phase cancellation (Cabrera, 2017). Von Neumann (1929) located entropy increase in the act of measurement, which introduces classical uncertainty. Modern approaches mostly make use of entanglement between a system and an environment to bring about increasing entropy, but, to make the issue meaningful, the system should be “quasi-isolated” (Lesovik et al., 2016). More precisely, the time scale is long enough to allow de-phasing of off-diagonal elements of the density matrix, but short enough to ensure negligible energy exchange with the environment (Lesovik et al., op.cit.).
In this section, we formulate an abstract H-theorem for signed measures. Our theorem can be thought of as encompassing, for example, the Wigner quasi-probability representation of quantum systems on phase space (Wigner, 1932) and other phase-space representations. (See Rundle and Everitt, 2021 for a recent survey.) More precisely, it will cover general evolution of any no-signaling system (Popescu and Rohrlich, 1994). This follows from Abramsky and Brandenburger (2011, Theorem 5.9), who prove an equivalence between no-signaling systems and signed phase-space representations. Also relevant here is van de Wetering (2018), who establishes quantum theory as a subcategory of the category of quasi-stochastic processes.
To proceed, fix again a finite set of states $X=\{x_{1},\ldots,x_{n}\}$, to be thought of as the set of possible phase states of a physical system. We allow for signed measures $P=(p_{1},\ldots,p_{n})$ on states in $X$. We next introduce transition rates $\lambda_{ij}$ between states $x_{i}$ and $x_{j}$, for $i,j=1,\ldots,n$. These rates are defined as limits:
${}\lambda_{ij}=\lim_{\Delta t\rightarrow 0}\frac{{\rm Pr}(x_{j}\,\,\text{at time}\,\,t+\Delta t\,|\,x_{i}\,\,\text{at time}\,\,t)-\delta_{ij}}{\Delta t},$ (24)
where $\delta_{ij}$ is the Dirac delta function: $\delta_{ij}=1$ if $i=j$ and 00 otherwise. As is customary, this limit is assumed to be well-defined (see, e.g., Whittle, 2000, p.183-184 for a discussion). Observe that $\sum_{j=1}^{n}\lambda_{ij}=0$ for each $i$, and, for a classical system, $\lambda_{ij}\geq 0$ for $i\not=j$.
To allow for fully general evolution of a signed measure on phase space, we first write down a master equation where $\lambda_{ij}<0$ is permitted. Call these signed transition rates. We impose micro-reversibility (Whittle, 2000, p.199): $\lambda_{ij}=\lambda_{ji}$ for all $i,j$. This assumption can be justified in various ways. In quantum mechanics, symmetry of transition rates is a consequence of Hermiticity of the Hamiltonian (Schwartz, 2021, Section 3.4). Our general master equation is:
$\frac{dp_{i}}{dt}$ $=\sum_{j=1}^{n}\lambda_{ji}p_{j}(t)$ (25)
$=\sum_{j=1}^{n}\lambda_{ji}p_{j}(t)-\sum_{j=1}^{n}\lambda_{ij}p_{i}(t)$ (26)
$=\sum_{j=1}^{n}\lambda_{ij}\bigl{[}p_{j}(t)-p_{i}(t)\bigr{]},~{}$ (27)
for $i=1,\ldots,n$. This is standard except that $P$ is a signed measure, and the matrix $\Lambda$ of transition rates may have negative off-diagonal entries.
The goal now is to identify a (minimal) condition on our master equation that yields a Second Law for signed Rényi entropy. The next theorem shows that such a condition is classicality of the evolution of the phase-space measure. Of course, this does not say that that phase-space measures must be non-negative, only that their evolution is ruled by non-negative probabilities. We state our result and then add some interpretation.
For our result, we consider a domain $\cal D$ in $\{P\in\mathbb{R}^{n}:w(P)\not=0\}$ that is closed under permutations, is convex, and has a non-empty interior. This will enable us to appeal to a differential characterization of Schur-concavity.
Theorem 4.
Assume the off-diagonal entries in the matrix $\Lambda$ of (signed) transition rates are non-negative – that is, assume classical evolution. Then signed Rényi $\alpha$-entropy, for $\alpha>1$, is non-decreasing on $\cal D$, that is: $dH^{\pm}_{\alpha}/dt\geq 0$.
Proof.
We find the time derivative of signed Rényi entropy:
${}\frac{dH^{\pm}_{\alpha}}{dt}=\sum_{i=1}^{n}\frac{\partial H^{\pm}_{\alpha}}{\partial p_{i}}\frac{dp_{i}}{dt}=\sum_{i=1}^{n}\frac{\partial H^{\pm}_{\alpha}}{\partial p_{i}}\sum_{j=1}^{n}\lambda_{ij}\bigl{[}p_{j}(t)-p_{i}(t)\bigr{]},$ (28)
where we have substituted in Equation 27. Interchanging $i$ and $j$, we can also write:
${}\frac{dH^{\pm}_{\alpha}}{dt}=\sum_{j=1}^{n}\frac{\partial H^{\pm}_{\alpha}}{\partial p_{j}}\frac{dp_{j}}{dt}=\sum_{j=1}^{n}\frac{\partial H^{\pm}_{\alpha}}{\partial p_{j}}\sum_{i=1}^{n}\lambda_{ji}\bigl{[}p_{i}(t)-p_{j}(t)\bigr{]}.$ (29)
Adding Equations 28 and 29 and using symmetry, we arrive at:
${}\frac{dH^{\pm}_{\alpha}}{dt}=-\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{n}\lambda_{ij}\bigl{(}\frac{\partial H^{\pm}_{\alpha}}{\partial p_{i}}-\frac{\partial H^{\pm}_{\alpha}}{\partial p_{j}}\bigr{)}\bigl{[}p_{i}(t)-p_{j}(t)\bigr{]}.$ (30)
By Theorem 3, signed Rényi entropy $H^{\pm}_{\alpha}$, for $\alpha>1$, is Schur-concave. Now, use the fact that a quotient of continuous functions with a strictly positive denominator is continuous, to see that $H^{\pm}_{\alpha}$ is continuous on $\cal D$. Next calculate:
$\frac{\partial H^{\pm}_{\alpha}}{\partial p_{j}}=-\frac{1}{(\alpha-1)\log_{e}2}\Bigl{[}\frac{\alpha|p_{j}|^{\alpha-1}\text{sign}(p_{j})}{\sum_{i}|p_{i}|^{\alpha}}-\frac{\text{sign}(\sum_{i}p_{i})}{|\sum_{i}p_{i}|}\Bigr{]}.$ (31)
The key observation here is that $\alpha|p_{j}|^{\alpha-1}\text{sign}(p_{j})$ is continuous on $\mathbb{R}$ precisely when $\alpha>1$, and the operations in Equation 31 preserve continuity. We conclude that $H^{\pm}_{\alpha}$ is $C^{1}$ on the interior of $\cal D$.
Using these properties, we can appeal to Marshall, Olkin, and Arnold (2011, Theorem A.4.a, pp.84-85) to obtain that for all $i,j$:
${}\bigl{(}\frac{\partial H^{\pm}_{\alpha}}{\partial p_{i}}-\frac{\partial H^{\pm}_{\alpha}}{\partial p_{j}}\bigr{)}\bigl{(}q_{i}-q_{j}\bigr{)}\leq 0,$ (32)
on $\cal D$. Since $\lambda_{ij}\geq 0$ for all $i\not=j$, it follows that $dH^{\pm}_{\alpha}/dt\geq 0$, as required. ∎
The key to Theorem 4 is the non-negativity of the transition rates. This can be thought of as a classicalization assumption, which, over time, has a smoothing effect on the system and thereby suppresses its ability to exhibit negative measure. We view our classicalization assumption as an abstraction of specific decoherence mechanisms posited for quantum systems in order to arrive at a quantum H-theorem. See the references at the beginning of this section, and also Gemmer, Michel, and Mahler (2009) and Gogolin and Eisert (2016) for comprehensive treatments.
Example 4.
We show that while signed Rényi entropy is non-decreasing, per Theorem 4, signed Shannon entropy may not be. Let the matrix of transition rates be given by:
$\Lambda=\begin{pmatrix}-1&1/2&1/2\\ 1/2&-1&1/2\\ 1/2&1/2&-1\end{pmatrix},$ (33)
and choose the initial signed probability measure: $P(0)=(-1/7,3/7,5/7)$. Figure 1 shows the evolution of signed Rényi $2$-entropy $H^{\pm}_{2}(t)$ and signed Shannon entropy $H^{\pm}_{1}(t)$. While signed Shannon entropy converges to the same limit $\log 3\approx 1.585$ as signed Rényi $2$-entropy, we see that it is transiently non-monotonic.
Figure 1: Evolution of Signed Rényi and Signed Shannon Entropies
7 Discrete Moyal-Bracket Evolution
Our H-Theorem hinged on non-negativity of off-diagonal entries of the matrix of transition rates – an abstraction of decoherence mechanisms investigated in the literature. An important check on this result is that signed Rényi entropy, for a suitably chosen free parameter, is constant and not increasing under discrete Moyal-bracket evolution on phase space. This will constitute the phase-space analog to the the standard fact that von Neumann entropy $H_{1}(\rho)=-\text{Tr}(\rho\log\rho)$ – and every Rényi entropy $H_{\alpha}(\rho)=-1/(\alpha-1)\times\log\text{Tr}(\rho^{\alpha})$ – is constant under unitary evolution on Hilbert space. The check will further underpin signed Rényi $\alpha$-entropy as a natural entropy for quasi-probability descriptions.
Fix the Hilbert-space dimension to be an odd-prime power and let $W$ denote the discrete Wigner function constructed from Wootters phase-point operators (Wootters, 1987) or, equivalently, from the finite field construction of Gibbons, Hoffman, and Wootters (2004). (The argument can be extended to general finite phase spaces using the methods in Ferrie, 2011.) Time evolution under a Hamiltonian $\cal H$ in this phase-space picture is governed by the discrete Moyal bracket (Gross, 2006):
$\dot{W}=\{{\cal H},W\}_{*d}.$ (34)
This evolution is implemented by a real skew-symmetric Liouvillian $L_{\cal H}$ acting on $W$:
$\dot{W}=L_{\cal H}W,\,\,L^{\mathsf{T}}_{\cal H}=-L_{\cal H}.$ (35)
Skew-symmetry holds because the discrete Moyal bracket inherits the antisymmetry of its continuous counterpart (Moyal, 1949).
Theorem 5.
Fix a discrete Wigner function $W$ and a Hamiltonian $\cal H$. Then $\sum_{\xi}W_{\xi}^{2}$ is a constant of motion, so that signed Rényi $2$-entropy is conserved, that is: $dH^{\pm}_{2}/dt=0$.
Proof.
Define the quadratic moment:
$V(t)=\sum_{\xi}W_{\xi}(t)^{2}=W(t)^{\mathsf{T}}\,W(t).$ (36)
Taking the time derivative:
$\frac{dV}{dt}=\bigl{(}\frac{dW^{\mathsf{T}}}{dt}\bigr{)}\,W\,+\,W^{\mathsf{T}}\bigl{(}\frac{dW}{dt}\bigr{)}=W^{\mathsf{T}}L_{\cal H}^{\mathsf{T}}\,W\,+\,W^{\mathsf{T}}L_{\cal H}W,$ (37)
where we interchange transpose and derivative and substitute in the first part of Equation 35. Substituting in from the second part:
$\frac{dV}{dt}=W^{\mathsf{T}}(-L_{\cal H})\,W\,+\,W^{\mathsf{T}}L_{\cal H}W=0,$ (38)
so that $V(t)$ is a constant of motion. It follows that:
$\frac{d}{dt}H^{\pm}_{2}\bigl{(}W(t)\bigr{)}=-\frac{d}{dt}\log V(t)=0,$ (39)
as required. ∎
8 Renormalization and Inverse Temperature
In this section we note a renormalization formula for the case $\alpha=1$ and also comment briefly on interpreting $\alpha$ as a (dimensionless) inverse temperature. First, we can easily renormalize our signed Rényi entropy formula $H^{\pm}_{\alpha}(P)$ at $\alpha=1$ by multiplying through by $\alpha-1$ to obtain:
$H^{\pm}_{1}(P)_{\text{Ren}}=-\log\bigl{[}\frac{\sum_{i=1}^{n}|p_{i}|}{|\sum_{i=1}^{n}p_{i}|}\bigr{]}.$ (40)
This formula satisfies Extensivity, witnesses negativity ($\sum_{i}|p_{i}|>|\sum_{i}p_{i}|$ if and only if at least one $p_{i}$ is negative), and is Schur-concave. However, it will generally fail to be conserved under Moyal-bracket evolution and so cannot serve as a fundamental entropy in a theory whose elementary reversible dynamics are unitary.
Next, we interpret the signed Rényi index $\alpha$ as a dimensionless inverse temperature parameter by writing $\alpha=T_{0}/T$, where we set the Boltzmann constant $k_{B}=1$ and $T_{0}$ is a fixed reference temperature.
Example 5.
For ordinary probability measures, Rényi entropy is well known to be non-increasing in $\alpha$. To examine the signed case, consider the signed probability measure $P=(\frac{4}{7},-\frac{1}{7},\frac{3}{14},\frac{5}{14})$. Figure 2 shows the behavior of signed Rényi entropy $H^{\pm}_{\alpha}(P)$ as a function of $1/\alpha$ over a range that produces nonmonotonicity. (We chose a range that does not include the pole at $\alpha=1$, since this is just a “reference-point nonmonotonicity” independent of any system structure.)
Figure 2: Nonmonotonicity of Signed Rényi Entropy
Interpreting $1/\alpha$ as temperature, we see that signed Rényi entropy first increases with temperature over the range considered, reaches a peak, and then decreases. While this example is just a toy model, it can be connected conceptually to quantum systems that display intervals of negative specific heat (Defenu et al., 2023). To the left of the peak, we are in the normal regime where increasing temperature goes with increasing entropy. Beyond the peak, we are in an “anomalous” negative specific-heat regime, where further heating decreases the entropy. (We note, though, that for a fixed signed probability measure, $H^{\pm}_{\alpha}(P)$ can display at most one extremum on either side of the pole. Obtaining an S-shaped caloric curve requires that the measure itself evolve; see Defenu et al., op.cit..)
9 Conclusion
We have provided a characterization of entropy for signed measures, arriving at the unique one-parameter family:
$H^{\pm}_{\alpha}(P)=-\frac{1}{\alpha-1}\log\bigl{[}\frac{\sum_{i=1}^{n}|p_{i}|^{\alpha}}{|\sum_{i=1}^{n}p_{i}|}\bigr{]},\quad\alpha>0\,\,\text{and}\,\,\alpha\not=1,$ (41)
singled out by our five information-theoretically interpretable axioms. The “obvious” signed Shannon entropy is excluded.
The operational virtues of our formula include: witnessing negativity, which is well understood as a key quantum resource (Veitch et al., 2012; Chabaud, Emeriau, and Grosshans, 2021); Schur-concavity, i.e., monotonicity under majorization for all $\alpha>1$, placing $H^{\pm}_{\alpha}$ in the family of entropy-like quantities that cannot decrease under mixing; again for $\alpha>1$, a quantum H-theorem for classicalizing dynamics; and, for $\alpha=2$, conservation under discrete Moyal-bracket (unitary) evolution. We also commented briefly on the physical meaning of the fact that signed Rényi entropy is non-monotonic in the $\alpha$ parameter.
An open foundational question is the extension of our axiomatic approach to obtain a useful formula for infinite phase space, complementing existing continuous entropy measures (Albarelli et al., 2018).
Appendix
In the section we prove Theorem 1. We start with the sufficiency direction and proceed via a number of lemmas.
Lemma A.1.
Under Axioms 0, 2, 3, and 4, if $p\not=0$, then $H((p))=-\log|p|$.
Proof.
Write $h(p)=H((p))$. Axioms 0 and 2 imply that $h$ is real-valued and continuous except at 00. Axiom 4 implies that $h(pq)=h(p)+h(q)$ whenever $p,q\not=0$. This is a version of Cauchy’s logarithmic functional equation, with general solution $h(p)=c\log|p|$, where $c$ is a real constant. See Aczél and Dhombres (1989, Theorem 3, p.27). Our Axiom 3 fixes $c=-1$. ∎
In fact, to obtain $h(q)=-\log|p|$, we can assume less than Axiom 2. See Corollary 2 and the development on pp.26-27 of Aczél and Dhombres (1989).
Lemma A.2.
Under Lemma A.1 and Axioms 4 and 5, we have $g(x)=-dx+e$ (linear) or $g(x)=d2^{(1-\alpha)x}+e$ (exponential), where $d\not=0$, $e$, and $\alpha\not=1$ are arbitrary constants.
Proof.
We adapt the argument in Daróczy (1963) to signed measures $Q$. From Lemma A.1 and induction on Axiom 5, we obtain:
${}H(P)=H((p_{1})\cup\cdots\cup(p_{n}))=g^{-1}\big{[}\frac{\sum_{j}w((p_{j}))g(H((p_{j})))}{w((p_{1})\cup\cdots\cup(p_{n}))}\big{]}=g^{-1}\big{[}\frac{\sum_{j}|p_{j}|g(-\log|p_{j}|)}{|\sum_{j}p_{j}|}\big{]}.$ (42)
From this and Axiom 4, we have for signed measures $P$ and $Q$, provided $\sum_{i,j}p_{i}q_{j}\not=0$:
${}g^{-1}\big{[}\frac{\sum_{i,j}|p_{i}q_{j}|g(-\log|p_{i}q_{j}|)}{|\sum_{i,j}p_{i}q_{j}|}\big{]}=g^{-1}\big{[}\frac{\sum_{i}|p_{i}|g(-\log|p_{i}|)}{|\sum_{i}p_{i}|}\big{]}+g^{-1}\big{[}\frac{\sum_{j}|q_{j}|g(-\log|q_{j}|)}{|\sum_{j}q_{j}|}\big{]}.$ (43)
Define $f:\mathbb{R}_{++}\rightarrow\mathbb{R}$ by $f(t)=g(-\log_{2}t)$. Substituting, we get:
${}f^{-1}\big{[}\frac{\sum_{i,j}|p_{i}q_{j}|f(|p_{i}q_{j}|)}{|\sum_{i,j}p_{i}q_{j}|}\big{]}=f^{-1}\big{[}\frac{\sum_{i}|p_{i}|f(|p_{i}|)}{|\sum_{i}p_{i}|}\big{]}\times f^{-1}\big{[}\frac{\sum_{j}|q_{j}|f(|q_{j}|)}{|\sum_{j}q_{j}|}\big{]}.$ (44)
Setting $Q=(q)$ (where $q\not=0$), this becomes:
${}\frac{1}{|q|}f^{-1}\big{[}\frac{\sum_{i}|p_{i}|f(|p_{i}q|)}{|\sum_{i}p_{i}|}\big{]}=f^{-1}\big{[}\frac{\sum_{i}|p_{i}|f(|p_{i}|)}{|\sum_{i}p_{i}|}\big{]}.$ (45)
Define $h_{q}:\mathbb{R}_{++}\rightarrow\mathbb{R}$ by $h_{q}(t)=f(|q|t)$. Then:
${}h_{q}^{-1}\big{[}\frac{\sum_{i}|p_{i}|h_{q}(|p_{i}|)}{|\sum_{i}p_{i}|}\big{]}=f^{-1}\big{[}\frac{\sum_{i}|p_{i}|f(|p_{i}|)}{|\sum_{i}p_{i}|}\big{]}.$ (46)
This establishes that the maps $h_{q}$ and $f$ generate the same means when restricting the $p_{i}$’s to be non-negative. By the necessity direction of a theorem on mean values (Hardy, Littlewood, and Pólya, 1952, Theorem 83), this implies that:
${}h_{q}(t)=a(q)f(t)+b(q),$ (47)
where $a(q)$ and $b(q)$ are independent of $t$, and $a(q)\not=0$. Substituting, we get:
${}f(|q|t)=a(q)f(t)+b(q).$ (48)
We return to follow Daróczy (1963) exactly and conclude that (restricting $q$ to be non-negative) the solution to Equation 48 takes the form:
${}f(t)=d\log_{2}t+e,$ (49)
or
${}f(t)=dt^{\alpha-1}+e,$ (50)
where $d\not=0$, $e$, and $\alpha\not=1$ are arbitrary constants. Recalling the definition of $f$, we then find that:
${}g(x)=-dx+e,$ (51)
or
${}g(x)=d2^{(1-\alpha)x}+e,$ (52)
as required. ∎
Lemma A.3.
Under Axiom 4, the function $g$ cannot be linear.
Proof.
If $g$ is linear as in Equation 51, then using Equation 42 we get:
${}-d\cdot H(P)+e=d\cdot\frac{\sum_{i}|q_{i}|\log|p_{i}|}{|\sum_{i}p_{i}|}+e\cdot\frac{\sum_{i}|p_{i}|}{|\sum_{i}p_{i}|},$ (53)
from which:
$H(P)=-\frac{\sum_{i}|p_{i}|\log|p_{i}|}{|\sum_{i}p_{i}|}-\frac{e}{d}\bigl{(}\frac{\sum_{i}|p_{i}|}{|\sum_{i}p_{i}|}-1\bigr{)}.$ (54)
Now let $P=(\frac{1}{2},\frac{1}{2})$ and $Q=(2,-1)$, so that $P*Q=(1,-\frac{1}{2},1,-\frac{1}{2})$. We calculate:
$H(P)=1,\,\,H(Q)=-2-\frac{2e}{d},\,\,H(P*Q)=1-\frac{2e}{d},$ (55)
so that:
$H(P*Q)\not=H(P)+H(Q),$ (56)
violating Axiom 4. ∎
Lemma A.4.
Under Axioms 0 and 4, the function $g$ must be exponential with $e=0$ and $\alpha>0$.
Proof.
If $g$ is exponential as in Equation 52, then again using Equation 42 we get:
${}d\cdot 2^{(1-\alpha)H(P)}+e=d\cdot\frac{\sum_{i}|p_{i}|^{\alpha}}{|\sum_{i}p_{i}|}+e\cdot\frac{\sum_{i}|p_{i}|}{|\sum_{i}p_{i}|},$ (57)
from which:
$H(P)=-\frac{1}{\alpha-1}\log\bigl{[}\frac{\sum_{i=1}^{n}|p_{i}|^{\alpha}}{\left|\sum_{i=1}^{n}p_{i}\right|}+\frac{e}{d}\bigl{(}\frac{\sum_{i}|p_{i}|}{|\sum_{i}p_{i}|}-1\bigr{)}\bigr{]}.$ (58)
Again let $P=(\frac{1}{2},\frac{1}{2})$ and $Q=(2,-1)$, so that $P*Q=(1,-\frac{1}{2},1,-\frac{1}{2})$. We calculate:
$H(P)=\log 2,\,\,H(Q)=-\frac{1}{\alpha-1}\log\bigl{[}2^{\alpha}+1+\frac{2e}{d}\bigr{]},\,\,H(P*Q)=-\frac{1}{\alpha-1}\log\bigl{[}2+2^{1-\alpha}+\frac{2e}{d}\bigr{]}.$ (59)
From Axiom 4:
$2^{1-\alpha}\times\bigl{[}2^{\alpha}+1+\frac{2e}{d}\bigr{]}=\bigl{[}2+2^{1-\alpha}+\frac{2e}{d}\bigr{]},$ (60)
which, for any $\alpha\not=1$, implies $e/d=0$.
Finally, if $\alpha<0$, then $H((0,1))$ is unbounded (negative), violating real-valuedness (Axiom 0). Therefore $\alpha\geq 0$. If $\alpha=0$, then $H(P)=1$ for all $Q$, violating non-constancy (Axiom 0). ∎
This completes the sufficiency direction of Theorem 1. To verify necessity, substitute $H^{\pm}_{\alpha}$ into Axioms 0, 2, 3, 4, and 5, and confirm each holds.